# Merging RLBWTs adaptively

### v4

@@ 1 Introduction
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts. In Section 3 we present our algorithm and in Section 4 we present a first analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 5 we give the proof of a technical lemma we use, whose details we refer to in Section 6 when we optimize our algorithm slightly to reduce the $r\log r$ in our time bound to $r$, so the whole bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), suffix arrays (SAs), the $\Psi$ function and longest common prefix (LCP) arrays; see Navarro’s [16] text for an introduction. The extended BWT (eBWT) [15] of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in the BWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$ of $S$ and $T$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
@@ 3 Main algorithm
(line 9) and reset $i$ to $i+k$ (line 10) and $j$ to $j^{\prime}$ (line 11). Since the characters we output are all equal, we need not consider their order. The case when there exists a character $\mathrm{BWT}_{T}[j+\ell]$ after the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ and $\mathrm{context}_{T}(j+\ell)\prec\mathrm{context}_{S}(i+k-1)$ (lines 13–16) is symmetric.
By the definition of the eBWT, it is impossible for $\mathrm{BWT}_{S}[i+k]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and for $\mathrm{BWT}_{T}[j+\ell]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$ in $\mathrm{BWT}_{S,T}$, since then $\mathrm{BWT}_{S}[i+k]$ would have to both precede and follow $\mathrm{BWT}_{T}[j+\ell]$. Therefore, the only remaining case with $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ is when either $i+k=m+1$ or $\mathrm{BWT}_{S}[i+k]$ follows the entire run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and either $j+\ell=n+1$ or $\mathrm{BWT}_{T}[j]$ follows the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$. In this case, we output

### v5

@@ 1 Introduction
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts. As a warm-up, in Section 3 we discuss an easy algorithm for merging positional BWTs (PBWTs). In Section 4 we present our main algorithm and in Section 5 we present a first analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 6 we give the proof of a technical lemma we use, whose details we refer to in Section 7 when we optimize our algorithm slightly to reduce the $r\log r$ in our time bound to $r$, so the whole bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), positional BWTs (PBWTs) [9], suffix arrays (SAs), the $\Psi$ function and longest common prefix (LCP) arrays; see Mäkinen et al.’s [17] and Navarro’s [19] texts for an introduction. The extended BWT (eBWT) [18] of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in the BWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$ of $S$ and $T$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
@@
3 Merging PBWTs
Suppose we have the PBWTs for two sets of haplotypes from the same species, with the same number of columns (but not necessarily the same number of rows) and with the columns representing the same variation sites, and we want the PBWT for all the haplotypes. For example, consider the PBWTs shown in Figure 1, with the bottom PBWT being the merge of the top two (and the example from [12] with indexing from 1 for consistency with the rest of this paper). The entries of the prefix arrays are shown in parentheses. Cells are red if their information is for one of the first 10 haplotypes and blue if it is for one of the second 10. This makes it easy to see that, by the definition of the PBWT, each column in the merged PBWT is an interleaving of the corresponding columns in the two input PBWTs.
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
1 (1) 1 (5) 0 (1) 1 (1) 1 (9) 0 (1) 0 (1) 0 (1) 0 (1) 0 (1) 0 (1) 1 (1) 1 (8) 1 (2) 1 (5)
1 (2) 1 (6) 0 (2) 1 (2) 0 (1) 1 (5) 0 (10) 0 (10) 1 (10) 1 (9) 0 (5) 1 (5) 0 (2) 1 (10) 1 (6)
1 (3) 1 (7) 0 (3) 1 (3) 1 (2) 1 (6) 0 (9) 0 (9) 0 (9) 0 (5) 0 (6) 1 (6) 0 (10) 0 (5) 1 (7)
1 (4) 1 (8) 0 (4) 1 (4) 1 (3) 1 (7) 0 (2) 1 (2) 0 (5) 0 (6) 0 (7) 1 (7) 1 (1) 0 (6) 1 (3)
0 (5) 1 (9) 0 (5) 1 (5) 1 (4) 1 (8) 0 (3) 1 (3) 0 (6) 0 (7) 0 (8) 0 (8) 0 (5) 0 (7) 1 (4)
0 (6) 1 (10) 0 (6) 1 (6) 0 (5) 0 (10) 0 (4) 1 (4) 0 (7) 0 (8) 0 (2) 0 (2) 0 (6) 0 (3) 1 (9)
0 (7) 0 (1) 0 (7) 1 (7) 0 (6) 0 (9) 0 (5) 0 (5) 0 (8) 0 (2) 0 (3) 1 (3) 0 (7) 0 (4) 1 (8)
0 (8) 0 (2) 0 (8) 1 (8) 0 (7) 0 (2) 0 (6) 0 (6) 0 (2) 0 (3) 0 (4) 1 (4) 0 (3) 0 (9) 1 (1)
0 (9) 0 (3) 0 (9) 0 (9) 0 (8) 0 (3) 0 (7) 0 (7) 0 (3) 0 (4) 0 (10) 0 (10) 0 (4) 0 (8) 1 (2)
0 (10) 0 (4) 0 (10) 1 (10) 0 (10) 0 (4) 0 (8) 0 (8) 0 (4) 0 (10) 1 (9) 1 (9) 0 (9) 0 (1) 1 (10)
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
0 (11) 1 (11) 0 (11) 1 (11) 1 (12) 0 (15) 0 (15) 0 (15) 1 (15) 0 (17) 0 (17) 1 (17) 1 (20) 1 (11) 1 (19)
0 (12) 1 (12) 0 (12) 0 (12) 1 (13) 0 (16) 0 (16) 0 (16) 1 (16) 0 (12) 1 (12) 1 (19) 1 (15) 0 (19) 0 (12)
0 (13) 1 (13) 0 (13) 0 (13) 1 (14) 1 (18) 0 (11) 0 (11) 1 (11) 0 (19) 0 (19) 1 (18) 1 (16) 0 (12) 1 (13)
0 (14) 1 (14) 0 (14) 0 (14) 0 (15) 0 (11) 0 (17) 0 (17) 0 (17) 0 (18) 0 (18) 0 (20) 0 (11) 0 (13) 1 (14)
0 (15) 1 (15) 0 (15) 0 (15) 0 (16) 0 (17) 0 (12) 0 (12) 0 (12) 0 (20) 0 (20) 0 (15) 1 (17) 0 (14) 1 (20)
0 (16) 1 (16) 0 (16) 0 (16) 0 (18) 0 (12) 0 (13) 0 (13) 1 (13) 0 (15) 0 (15) 0 (16) 0 (19) 0 (20) 1 (15)
0 (17) 1 (17) 0 (17) 1 (17) 1 (19) 0 (13) 0 (14) 0 (14) 1 (14) 0 (16) 0 (16) 0 (11) 1 (18) 0 (15) 1 (16)
1 (18) 1 (19) 1 (19) 0 (18) 1 (20) 0 (14) 0 (19) 0 (19) 0 (19) 0 (11) 0 (11) 1 (12) 0 (12) 0 (16) 1 (17)
0 (19) 1 (20) 1 (20) 0 (19) 0 (11) 0 (19) 1 (20) 0 (18) 0 (18) 0 (13) 1 (13) 1 (13) 0 (13) 0 (17) 1 (18)
0 (20) 1 (18) 0 (18) 0 (20) 0 (17) 0 (20) 0 (18) 0 (20) 0 (20) 0 (14) 1 (14) 1 (14) 0 (14) 0 (18) 1 (11)
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
1 (1) 1 (5) 0 (1) 1 (1) 1 (9) 0 (15) 0 (15) 0 (15) 1 (15) 0 (1) 0 (1) 1 (1) 1 (8) 1 (2) 1 (19)
1 (2) 1 (6) 0 (2) 1 (2) 1 (12) 0 (16) 0 (16) 0 (16) 1 (16) 0 (17) 0 (17) 1 (17) 1 (20) 1 (10) 1 (5)
1 (3) 1 (7) 0 (3) 1 (3) 1 (13) 1 (18) 0 (1) 0 (1) 0 (1) 1 (9) 1 (12) 1 (19) 0 (2) 1 (11) 1 (6)
1 (4) 1 (8) 0 (4) 1 (4) 1 (14) 0 (1) 0 (10) 0 (10) 1 (10) 0 (12) 0 (19) 1 (18) 1 (15) 0 (19) 1 (7)
0 (5) 1 (9) 0 (5) 1 (5) 0 (15) 1 (5) 0 (11) 0 (11) 1 (11) 0 (19) 0 (18) 1 (5) 1 (16) 0 (5) 1 (3)
0 (6) 1 (10) 0 (6) 1 (6) 0 (16) 1 (6) 0 (17) 0 (17) 0 (17) 0 (18) 0 (5) 1 (6) 0 (10) 0 (6) 1 (4)
0 (7) 1 (11) 0 (7) 1 (7) 0 (18) 1 (7) 0 (9) 0 (9) 0 (9) 0 (5) 0 (6) 1 (7) 0 (11) 0 (7) 0 (12)
0 (8) 1 (12) 0 (8) 1 (8) 1 (19) 1 (8) 0 (12) 0 (12) 0 (12) 0 (6) 0 (7) 0 (8) 1 (1) 0 (3) 1 (13)
0 (9) 1 (13) 0 (9) 0 (9) 1 (20) 0 (10) 0 (13) 0 (13) 1 (13) 0 (7) 0 (8) 0 (20) 1 (17) 0 (4) 1 (14)
0 (10) 1 (14) 0 (10) 1 (10) 0 (1) 0 (11) 0 (14) 0 (14) 1 (14) 0 (8) 0 (20) 0 (2) 0 (19) 0 (12) 1 (9)
0 (11) 1 (15) 0 (11) 1 (11) 1 (2) 0 (17) 0 (19) 0 (19) 0 (19) 0 (20) 0 (2) 1 (3) 1 (18) 0 (13) 1 (8)
0 (12) 1 (16) 0 (12) 0 (12) 1 (3) 0 (9) 1 (20) 1 (2) 0 (18) 0 (2) 0 (3) 1 (4) 0 (5) 0 (14) 1 (20)
0 (13) 1 (17) 0 (13) 0 (13) 1 (4) 0 (12) 0 (2) 1 (3) 0 (5) 0 (3) 0 (4) 0 (15) 0 (6) 0 (9) 1 (15)
0 (14) 1 (19) 0 (14) 0 (14) 0 (5) 0 (13) 0 (3) 1 (4) 0 (6) 0 (4) 0 (15) 0 (16) 0 (7) 0 (8) 1 (16)
0 (15) 1 (20) 0 (15) 0 (15) 0 (6) 0 (14) 0 (4) 0 (18) 0 (7) 0 (15) 0 (16) 0 (10) 0 (3) 0 (20) 1 (1)
0 (16) 0 (1) 0 (16) 0 (16) 0 (7) 0 (19) 0 (18) 0 (5) 0 (8) 0 (16) 0 (10) 0 (11) 0 (4) 0 (15) 1 (17)
0 (17) 0 (2) 0 (17) 1 (17) 0 (8) 0 (20) 0 (5) 0 (6) 0 (20) 0 (10) 0 (11) 1 (12) 0 (12) 0 (16) 1 (18)
1 (18) 0 (3) 1 (19) 0 (18) 0 (10) 0 (2) 0 (6) 0 (7) 0 (2) 0 (11) 1 (13) 1 (13) 0 (13) 0 (1) 1 (2)
0 (19) 0 (4) 1 (20) 0 (19) 0 (11) 0 (3) 0 (7) 0 (8) 0 (3) 0 (13) 1 (14) 1 (14) 0 (14) 0 (17) 1 (10)
0 (20) 1 (18) 0 (18) 0 (20) 0 (17) 0 (4) 0 (8) 0 (20) 0 (4) 0 (14) 1 (9) 1 (9) 0 (9) 0 (18) 1 (11)
Figure 1: Two PBWTs for 10 haplotypes each (above), and the PBWT for all 20 haplotypes (below). The entries of the prefix arrays are shown in parentheses. Cells are red if their information is for one of the first 10 haplotypes and blue if it is for one of the second 10.
The first column from the left of the merged PBWT is the concatenation of the the first columns of the two input PBWTs. Suppose we have already merged the (j1)(j-1)st columns of the input PBWTs into the (j1)(j-1)st column of the merged PBWT, and now we want to merge their $j$th columns into its $j$th column. In particular, suppose we know how the (j1)(j-1)st column of the merged PBWT is divided into maximal blocks of consecutive bits that all come from the same input PBWT. In our example, the blocks in the 9th column have lengths $2,2,2,1,5,4,1,3$ and alternate between blue and red with the first block being blue.
To compute the $j$th column of the merged PBWT after computing the (j1)(j-1)st column, we first consider all the 0s in the (j1)(j-1)st column and list the next bits in their haplotypes (in the 0s’ order in the (j1)(j-1)st column), then consider all the 1s in the (j1)(j-1)st column and list the next bits in their haplotypes (in the 1s’ order in the (j1)(j-1)st column); the $j$th column of the merged PBWT is the concatenation of the two lists. By induction, the bits in the $j$th column are in the co-lexicographic order of the preceding prefixes in their haplotypes.
Because the $j$th columns of the two input PBWTs already contain all those next bits and in the correct order, we can use the input PBWTs instead of the explicit haplotypes. To do this, we start with an empty draft of the $j$th column of the merged PBWT and scan the (j1)(j-1)st column of the merged PBWT twice. During the first scan, whenever we see a 0 we copy a bit (the next bit in that 0’s haplotype) from the $j$th column of the input PBWT that is the source of that 0. During the second scan, whenever we see a 1 we copy a bit (the next bit in that 1’s haplotype) from the $j$th column of the input PBWT that is the source of that 1. However, we may be able to work more efficiently considering blocks rather than literally bit by bit.
Suppose the (j1)(j-1)st column of the merged PBWT consists of $b$ blocks. For $k$ from 1 to $b$, if there are $c$ copies of 0 in the $k$th block in the (j1)(j-1)st column then we append to our draft of the $j$th column of the merged PBWT the next $c$ bits of the $j$th column of the input PBWT that is the source of the bits in the $k$th block. After that, for $k$ from 1 to $b$, if there are $c$ copies of 1 in the $k$th block of the (j1)(j-1)st column then we append to our draft of the $j$th column of the merged PBWT the next $c$ bits of the $j$th column of the input PBWT that is the source of the bits in the $k$th block.
In our example, there are no copies of 0s in the first block of the 9th column, which is blue, so we append no bits from the 10th column of the blue PBWT; there is 1 copy of 0 in the second block of the 9th column, which is red, so we append 1 bit (0) from the 10th column of the red PBWT (from haplotype (1)); then 1 bit (0) from the blue PBWT (from haplotype (17)); then 1 bit (1) from the red PBWT (from haplotype (9)); then 3 bits (000) from the blue PBWT (from haplotypes (12), (19) and (18)); then 4 bits (0000) from the red PBWT (from haplotypes (5), (6), (7) and (8)); then 1 bit (0) from the blue PBWT (from haplotype (20)); then 3 bits (000) from the red PBWT (from haplotypes (2), (3) and (4)).
There are 2 copies of 1 in the first block of the 9th column, which is blue, so we append 2 bits (00) from the 10th column of the blue PBWT (from haplotypes (15) and (16)); then 1 bit (0) from the from the red PBWT (from haplotype (10)); then 1 bit (0) from the blue PBWT (from haplotype (11)); then no bits from the red PBWT; then 2 bits (00) from the blue PBWT (from haplotypes (13) and (14)). When we are finished, we know the 10th column of the merged PBWT is 00100000000000000000 and we know its blocks have lengths 1, 1, 3, 4, 1, 3, 2, 1, 3; its prefix array is (1) (17) (9) (12) (19) (18) (5) (6) (7) (8) (20) (2) (3) (4) (15) (16) (10) (11) (13) (14).
If a block in the (j1)(j-1)st column of the merged PBWT overlaps $t$ runs of bits in that column, then computing the numbers of 0s and 1s in that block takes $O(t)$ time. We perform one append operation for the 0s in that block and another for the 1s. If the input PBWTs are run-length compressed and we want only the $j$th column and its block structure — which is all we need to continue merging the two input PBWTs — but not its prefix array, then each append operation takes time proportional to the number of runs in the substring of bits we append. It follows that we can merge two PBWTs in time $O(r+B)$, where $r$ and $B$ are the total numbers of runs and blocks in all the columns of the merged PBWT, respectively.
Theorem 3.1.
We can merge two PBWTs with run-length compressed columns in $O(r+B)$ time, where $r$ and $B$ are the total numbers of runs and blocks in all the columns of the merged PBWT, respectively.
Theorem 3.1 may be of independent interest but we have proven it mainly as a warm-up for merging RLBWTs, which is significantly more complicated.
@@ 4 Main algorithm
(line 9) and reset $i$ to $i+k$ (line 10) and $j$ to $j^{\prime}$ (line 11). Since the characters we output are all equal, we need not consider their order to merge the RLBWTs. (We note, however, that we would need to consider the characters’ order if we were also merging the RLBWTs’ SAs.) The case when there exists a character $\mathrm{BWT}_{T}[j+\ell]$ after the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ and $\mathrm{context}_{T}(j+\ell)\prec\mathrm{context}_{S}(i+k-1)$ (lines 13–16) is symmetric.
By the definition of the eBWT, it is impossible for $\mathrm{BWT}_{S}[i+k]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and for $\mathrm{BWT}_{T}[j+\ell]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$ in $\mathrm{BWT}_{S,T}$, since then $\mathrm{BWT}_{S}[i+k]$ would have to both precede and follow $\mathrm{BWT}_{T}[j+\ell]$. Therefore, the only remaining case with $\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]$ is when either $i+k=m+1$ or $\mathrm{BWT}_{S}[i+k]$ follows the entire run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and either $j+\ell=n+1$ or $\mathrm{BWT}_{T}[j]$ follows the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$. In this case, we output